Graph the line represented by 2x-3y=-6
Answer by benni1013(206)

Your line is in the standard form of an equation we need to turn it into the slope-intercept form.
2x+3y=-6==>3y=-6-2x==>y=-2-2/3x==>y=-2/3x-2
Plot the point at (0,-2) go down 2 over 3 plot point and draw line.